OpenStudy (anonymous):

If 7x+2y-6z=12 and if x, y, and z are positive, then what is the value of (2+z)/7x+2y ?

hartnn (hartnn):

can u isolate 7x+2y from 7x+2y-6z=12

OpenStudy (anonymous):

@hartnn Would it be 6/12?

hartnn (hartnn):

7x+2y-6z=12 add 6z on both sides, what u get ? ....wait 6/12 as final answer ?

OpenStudy (anonymous):

yep!

hartnn (hartnn):

correct! good work! :)

OpenStudy (anonymous):

@hartnn Wait, that's not part of the answer choices!

OpenStudy (anonymous):

Would it be 1/12?

hartnn (hartnn):

not 1/12 ... 6/12 can be simplified!

OpenStudy (anonymous):

I know, 1/2 is not on there either /:

hartnn (hartnn):

:O

hartnn (hartnn):

list the choices...

hartnn (hartnn):

7x+2y = 12+6z = 6 (2+z) (7x+2y)/(2+z) = 6 (2+z)/(7x+2y) = 1/6 sorry :\

OpenStudy (anonymous):

@hartnn No, I'm sorry! I keep making stupid mistakes

hartnn (hartnn):

we did same mistakes :P